### v1

# An iterated learning model of language change that mixes supervised and unsupervised learning

Abstract
The iterated learning model is an agent-based model of language change in which language is transmitted from a tutor to a pupil which itself becomes a tutor to a new pupil, and so on. Languages that are stable, expressive, and compositional arise spontaneously as a consequence of a language transmission bottleneck. Languages are expressive to the extent that a range of different meanings are conveyed unambiguously, and are compositional to the extent that the component parts of meanings are mapped consistently to component parts of the signals that convey them. In the model, a language-using tutor provides a naïve pupil with a limited training set of examples that each pair a meaning with the signal that conveys it in the tutor’s language. The transmission bottleneck ensures that pupils must generalize from the limited training set of meaning-signal pairs. After training, the pupil tutors a new naïve agent, and the cycle repeats. Under the right conditions, the repeated cycles of learning and generalization result in a language that is stable, expressive and compositional. Previous models have implemented an agent’s mapping from signals to meanings using an artificial neural network decoder, but have relied on an unrealistic and computationally expensive process of obversion to implement the associated encoder, mapping from meanings to signals. Here, a new model is presented in which both decoder and encoder are neural networks, trained separately through supervised learning, and trained together through unsupervised learning in the form of an autoencoder. This avoids the substantial computational burden entailed in obversion and introduces a mixture of supervised and unsupervised learning observed during human development. The model achieves more complex expressive compositional languages than previous studies, and reveals a linear relationship between the dimensionality of a language’s meaning-signal space and the size of the optimal transmission bottleneck.
Introduction
Languages are made easier to learn and easier to interpret by their expressivity, compositionality and stability; that is, by the tendency for a language to be able to express many different meanings using different sentences, to compose meaningful sentences from meaningful sub-parts and to change slowly over generations. There are different words for left and right, for example, but the word ‘left’ in ‘you write with your left hand’ means something very similar to the word ‘left’ in ‘turn left at the junction’. Moreover, the difference in meaning between the expressions ‘turn left, then right’ and ‘turn right, then left’ is represented by the different structural organization of their component words. Of course, languages are neither completely expressive nor completely compositional, ‘they are on the left’ typically refers to position but, in some contexts, could refer to politics, and a sailor in some cases would use the word ‘port’ when left might be expected. Nonetheless, expressivity and compositionality are, invariably, important features of languages and might be considered to mark the boundary between a simple system of signs and utterances and a true language. The third property, stability, is also important. Since parents need to be able to understand what their children say to them, at least some of the time, it is important that language does not change drastically from generation to generation. Obviously these are only three of a long list of properties a ‘good’ language needs in order that it be easy to learn and effective for communication; however, these three are among the most basic and are our focus here.
The iterated learning model
The iterated learning model (ILM) is an agent-based model of language change that demonstrates how a transmission bottleneck can encourage the spontaneous emergence of a language that is expressive, compositional and stable. Here a language is defined as a map from a set of ‘meanings’ to a set of ‘signals’ and, broadly, the goal for the iterated learning model is to shed light on the conditions under which languages that are expressive, compositional and stable arise. We will define measures, $x$, $c$ and $s$ for expressivity, compositionality and stability; these measures all take values in $[0,1]$ and we define an XCS language as one where all three exceed $\lambda=0.95$. In our limited context this will serve as a definition of a ‘good’ language.
The ILM was introduced in [1]. In its original form, it demonstrated that the emergence of language compositionality could be attributed to the influence of a transmission bottleneck on language acquisition taking place over a number of successive generations [1, 2]. However, this original version of the ILM did not include neural networks. Instead, languages were modeled using simple formal grammars. The version of the ILM considered here, with a decoder based on a neural network, appeared in [3]. This incorporated obversion, an algorithm to derive an encoding map from the decoding map, which had been previously defined in a more general context in [4]. Since then, obverter based learning models have been applied to pictures [5] and population-based versions of the ILM have been considered [6, 7, 8]. The ILM has also been discussed in a broader Bayesian framework, [9, 10] and this has been applied to populations [11] and to cultural evolution [12, 13]. Similar models of language development include the talking heads experiment, [14], the seeded ILM [15], the neural iterated learning model, [16], and an extension of what is called the restaurant process, [17], to language [18]. These modeling-based approaches also have a counterpart in participant studies using artificial mini-languages, where similar issues of expressivity, compositionality and stability are also considered, see [19, 20, 21] for reviews.
In the ILM, agents can decode signals, that is, map them to meanings, and can encode meanings, that is map them to signals:
$\mathcal{S}$$\mathcal{M}$$e$$d$
Here, $\mathcal{S}$ and $\mathcal{M}$ are the sets of all possible signals and all possible meanings, respectively, $e$ is a particular encoder and $d$ is a particular decoder.
A tutor agent with its own private encoder and decoder mappings is charged with teaching its language to an initially naïve pupil agent, that is, an agent with initial mappings that are empty, random or arbitrary. To do this, the tutor presents a series of meaning-signal pairs to the pupil, where each signal is created by encoding a target meaning using the tutor’s encoder. The presentation of each example is an instance of supervised learning, like an adult pointing to a boy stealing biscuits from a cookie jar and saying ‘Look! The boy is stealing biscuits from the cookie jar’. The pupil uses these examples to train its own decoder, learning a mapping between signals and meanings. After this training, the pupil ‘grows up’ and becomes a tutor itself. To act as a tutor it also requires its own encoder, which is derived from its trained decoder by a process of ‘obversion’. The new tutor next trains a new naïve pupil who again initially starts with language mappings that are empty, random or arbitrary, and in this way the process iterates.
Despite each pupil starting its training in the same naïve state with no knowledge of the language that the tutor has developed, the ILM demonstrates a tendency for the language that is being learned to change systematically over successive agent ‘lifetimes’. Under certain conditions the language stabilizes so that successive agents learn the same language and this language exhibits high levels of expressivity and compositionality.
Whether or not the ILM achieves this kind of outcome depends on the size of its transmission bottleneck: the number of meaning-signal pairs that the tutor presents to the pupil during training. Since this is a subset of the full set of meaning-signal pairs, when the pupil becomes a tutor they must generalize from what they have learned in order to be able to present their pupil with a new set of meaning-signal pairs comprising random meanings, each paired with the signal that encodes it in the tutor’s language. Where the bottleneck is too small, language does not stabilize. Where the bottleneck is too large, language may stabilize on an expressive but non-compositional language, that is, one in which the relationship between each meaning and its associated signal is arbitrary and unstructured.
The neural-network based ILM
Here, the neural-network based ILM presented in [3] will be referred to as the Obverter ILM to distinguish it from the Semi-Supervised ILM presented in this paper. In the Obverter ILM, and indeed in all the ILMs presented in this paper, the meanings and signals are both binary $n$-vectors so that $|\mathcal{M}|=|\mathcal{S}|=2^{n}$. An individual bit, $i$, in a meaning vector $m\in\mathcal{M}$ will be called a ‘fact’. A bit, $j$, in a signal vector $s\in\mathcal{S}$ will be called a ‘word’, in fact, ‘morpheme’ would be more accurate, since the signal vector is intended to model any aspect of how meaning is encoded, but, for convenience, ‘word’ will be used. An agent has two maps, an encoding map from meaning to signal:
$e:\mathcal{M}\rightarrow\mathcal{S}$ (1)
and a decoding map from signal to meaning:
$d:\mathcal{S}\rightarrow\mathcal{M}$ (2)
A - Words and facts B - A language    C - A decoder
[…]
Fig 1: A guide to the notation. A: A meaning is an ordered sequence of $n$ facts, $m_{1}$ thru $m_{n}$. A signal is an ordered sequence of $n$ words, $s_{1}$ thru $s_{n}$. The use of word is potentially confusing since a word is sometimes thought of as a sequence of letters, but here it is an indivisible component of the signal. A word corresponds to a single bit and a signal can be thought of as corresponding to a phrase; in the same way a meaning can be thought of as corresponding to a state of the world, composed of a set of facts. B: This is an example $n=3$ language; it is fully expressive: every possible meaning maps to a different signal and it is fully compositional: each fact fully determines a unique word. Here, $s_{1}=\neg m_{3}$, $s_{2}=m_{1}$ and $s_{3}=\neg m_{2}$. For illustrative convenience the $m_{2}$ and $s_{3}$ elements involved in the last of these equivalences have been colored orange. In the example in C the decoder map $d$ maps the signal $(1,0,\ldots,1)$ to the meaning $(0,1,\ldots,1)$; the decoder is made up of two parts, the neural network $\hat{d}$ with one hidden layer the same size as the input and output, and the decider map $\delta$ which maps probabilities to zero or one.
The decoder, $d:\mathcal{S}\rightarrow\mathcal{M}$, is a neural network with input, hidden and output layers each of size $n$. It is useful to be careful at this point with the notation; the neural network itself maps a meaning, $m\in\mathcal{M}$, to a vector $(p_{1},p_{2},\ldots,p_{n})$ where $p_{i}\in[0,1]$ is the probability that the $i$th word in the corresponding signal is one. This map is called $\hat{d}$:
$\hat{d}:\mathcal{M}\rightarrow\mathcal{P}_{n}$ (3)
where $\mathcal{P}_{n}$ is the space of $n$-vectors whose entries are all probabilities. To add further to the notation, a decision map $\delta$ is defined, this maps a vector of probabilities to a binary vector by mapping a probability $p>0.5$ to one and $p<0.5$ to zero. In this way $d(m)=\delta(\hat{d}(m))$, or for those who prefer more formal notation: $d=\delta\circ\hat{d}$. As a final note, $\hat{d}$ also defines a map from a meaning and signal pair to a probability:
$\hat{d}(m,s)=\prod p(m_{i})$ (4)
where $\hat{d}(m)=(p_{1},p_{2},\ldots,p_{n})$ and $p(m_{i})=p_{i}$ if $m_{i}$ is one and $1-p_{i}$ if $m_{i}$ is zero. There is a guide to the notation in Fig. 1.
The encoder, $e$, is more complicated; it takes the form of a list tabulating the preferred signal corresponding to each meaning in $\mathcal{M}$. This list is calculated by obversion, an inversion-like calculation performed on the decoder map when an agent transitions from pupil to tutor. Obversion is described in more detail in S1 Appendix., but the key point is that it requires the calculation of the $2^{2n}$ entries of the table of probabilities given by $\hat{d}(m,s)$ for all $(m,s)$ pairs. This is costly, already for $n=8$ the table of probabilities is cumbersome, having 65,536 entries. As $n$ grows it rapidly becomes too large to calculate easily. It is also unrealistic. One of the signature properties of language is its open-ended power to encode novel meanings. It is certainly hard to imagine that when a child is learning language they conjure up all possible subjects of future discourse and estimate the probability that each would be expressed by every possible sentence. The main aim of this paper is to introduce a new ILM which does not require obversion and to confirm that language expressivity, compositionality and stability still arise under conditions analogous to those required for the original Obverter ILM.
[…]
Fig 2: Training the Obverter ILM. Each agent, $A_{i}$, first trains its decoder during a period of supervised learning on a set of meaning-signal pairs, ${{\cal{B}}_{i}}$, provided by its tutor, $A_{i-1}$. Subsequently, the pupil (red) derives an encoder from its decoder using a process of obversion and itself is promoted to become a tutor (blue) to a new pupil agent, $A_{i+1}$.
[…]
Fig 3: Training the Semi-Supervised ILM. Each agent, $A_{i}$, trains their encoder and decoder during a period combining supervised learning on a set of meaning-signal pairs, $\mathcal{B}_{i}$, provided by its tutor $A_{i+1}$ and unsupervised learning on a set of example meanings, $\mathcal{A}_{i}$ drawn from $\mathcal{M}$. Subsequently, the pupil (red) is promoted to become a tutor (blue) to a new pupil agent, $A_{i+1}$; unlike in the Obverter ILM, this promotion does not involve any further work for the agent since its encoder is already trained.
In the Obverter ILM the pupil starts as a naïve agent; it is presented with examples of signal-meaning pairs by the tutor agent and uses these to train its decoder map, $\hat{d}$, using stochastic gradient descent. The set of example meaning-signal pairs, $\mathcal{B}$, is produced by randomly selecting a subset of meanings and mapping these to signals using $e$, the tutor’s encoder. The set $\mathcal{B}$ is smaller than the set of meanings and so it constitutes a bottleneck in language transmission: in particular the set of examples a tutor presents will generally be different from the examples they learned from and so the tutor is forced to generalize beyond its training corpus. This is an important element of the ILM and the set $\mathcal{B}$ is referred to as the bottleneck set. After training, the pupil obverts its decoder map, $d$, to construct its encoder map, $e$, and becomes the tutor to a new naïve pupil. This cycle is repeated many times in order to explore how language structure changes over multiple generations. Fig. 2 gives a diagram of the generational structure.
A new ILM
In this paper we propose a novel Semi-Supervised ILM that employs a different agent architecture; in addition to using a neural network for the decoding map, mapping signals to meanings, it also uses a separate neural network for the encoding map, mapping meanings to signals, and, crucially, an autoencoder, mapping meanings through signals and back to meanings. This is described in detail below, but, roughly speaking, the autoencoder is formed by concatenating the decoder and encoder networks. During training, the pupil is sometimes given a meaning-signal pair as part of a supervised learning trial on either its encoder or decoder, but sometimes it is just given a meaning as an unsupervised learning trial on the autoencoder network as a whole. This unsupervised learning might be considered akin to a child observing what is happening around them and imagining how it might be described. This additional element of the training appears crucial to the development of an expressive, compositional and stable language. Fig. 3 gives a diagram of this generational structure.
Dispensing with the ILM’s obverter improves its ecological validity, that is, it moves the model closer to the situation it is intended to shed light on. The particular semi-supervised approach employed here seems especially apt for modeling language learning since human language learning does include a mix of explicit instruction and implicit exposure. Furthermore, as a result of escaping the computational costs of obversion, this version of the ILM supports larger ‘languages’. This should prove useful when using the ILM to study language contact or language change in structured communities [6, 7, 8] and should also make it easier to incorporate and study other aspects of language structure and development.
Results
The Obverter ILM
The performance of the Obverter ILM is reprised in Fig. 4. As in [3] it is demonstrated that an XCS language can quickly arise within the ILM: Fig. 4A-C show that after 40 generations all three properties are close to their maximum. These graphs are all based on a bottleneck size of 50 meaning-signal pairs, which is shown to be an approximately optimal size in Fig. 4D-F.
A B C
Refer to caption Refer to caption Refer to caption
D E F
Refer to caption Refer to caption Refer to caption
Fig 4: The Obverter ILM evolves a stable, expressive, compositional language. A-C plot expressivity ($x$), compositionality ($c$) and stability ($s$) as a function of generation for 25 independent replicates of the $n=8$ Obverter ILM; the thick line is the mean, the thin lines show the individual replicates and the bottleneck size in all cases is 50. D-F plot mean values for $x$, $c$ and $s$ as a function of bottleneck size after 40 generations (solid line) and 15 generations (dashed line) for 25 independent replicates per data point. The ribbons depict standard deviations around the 40-generation mean.
Figure 5 is perhaps more surprising. This set of simulations examines the ideal size of the bottleneck as $n$, the bit-length of the meaning and signal spaces, is changed. For this a precise, though of course slightly arbitrary, definition of ideal size is used: the model is run until $x$, $c$ and $s$ all exceed a $\lambda=0.95$ threshold, indicating that an XCS language has developed. The number of generations is noted, this is repeated 25 times and the mean calculated. The size of the bottleneck which enables the threshold to be reached most quickly, on average, is taken as the best bottleneck size. It turns out that this value increases linearly with $n$. This is despite the size of the meaning and signal sets increasing exponentially with $n$; for $n=8$ the optimal bottleneck is close to a quarter the number of meanings, while for $n=12$ it is only about a 40th.
A B
Refer to caption Refer to caption
Fig 5: The ideal bottleneck size is smaller than you might expect. For each bottleneck size considered, 25 replicates of the model are run until values of an XCS language develops. The best value of the bottleneck was estimated and this is plotted in A; the line shows the standard best linear fit, with slope 8.4 and intercept -11.5. The mean number of generations taken to reach the $\lambda$ threshold at this bottleneck size is plotted in B.
Perhaps the simplest approach to avoiding obversion is to model only one half of the language learning task and use agents with only an encoder. In this approach the tutor is given a set of meanings to present to the pupil and uses its encoder to pair them with associated signals. The pupil uses this set of meaning-signal pairs to train its own encoder, and then becomes a tutor for a new naïve pupil. This approach is obviously less satisfactory than the full ILM because it does not consider the ability of agents to decode signals into meanings. Moreover, it does not work; this one-way ILM always fails to evolve an expressive language. On the contrary, a stable language in which most if not all meanings are mapped to the same signal or a small set of similar signals arises as this type of mapping is readily learnable from a small set of training examples and is stable in the absence of any pressure to avoid ambiguity. This result also points to an effect of the obversion process beyond acting as a convenient approach to constructing $e$ from $\hat{d}$: by picking a ‘winner’ among the signals proposed to correspond to a given meaning, the obverter encourages expressivity, though it does not in any formal sense, guarantee that the encoder is expressive.
The Semi-Supervised ILM
It is possible that the collapse of expressivity exhibited by the one-way ILM could be prevented by adding a contrastive term to the model’s objective function. In this approach the neural network encoder is rewarded, as usual, for mapping a signal to the same meaning as the tutor supplied. However, in addition, the network would be punished if it allowed two recently witnessed signals to be mapped to the same or similar meanings. Our own experiments with this approach have not been successful and our belief is that this is because the contrastive term only deals with the meaning-signal pairs in the bottleneck set: the collapse in expressivity seems to be result of generalisation tending to associate unseen meanings with signals that were experienced during learning. In the remainder of this paper a different approach is taken.
In the Semi-Supervised ILM, a new agent architecture is considered which uses one neural network for the decoder and another for the encoder. In the Obverter ILM the bottleneck set is formed by randomly selecting a subset of meanings and then matching them to signals using the tutor’s encoder. In the Semi-Supervised ILM both the pupil’s encoder and decoder need to be trained so example meaning-signal pairs from $\mathcal{B}$ are used as meaning-signal pairs to train the encoder and as signal-meaning pairs to train the decoder. It is easy to see, however, that this simple approach will not work on its own. Since the decoder does not contribute to constructing $\mathcal{B}$ it has no role in the evolution of the language and is, as it were, ‘just along for the ride’. Consequently, while the addition of the decoder network may make the agent more realistic, from a language learning point-of-view it serves no function and, as far as the language is concerned, this ILM is no different from the one-way ILM described above.
To resolve this it is useful to consider the situation with human language. During human language learning, only a portion of a child’s language exposure is ‘supervised’ in the sense that an adult purposefully and explicitly presents a signal-meaning pair. It is clear, on reflection, that the presentation of learning examples to a child is complex, with some ‘supervised’ child-directed speech including clear indications of meaning, some wholly ‘unsupervised’ learning based on overheard speech or based on cogitation during observation, and a range of intermediate examples where a child hears utterances related to some activity of situation that they are engaged in, or where some indication of meaning is available or can be inferred, but has not been made explicit or precise.
This is an important and well-studied topic, see[22, 23, 24, 25, 26, 27] for example. It is also a complex topic, with child-adult interactions changing as the child becomes older and without any culturally-preserved norms around the structure of a child’s language exposure. Crucially, a classic study, [28] replicated in [29, 30], demonstrates that during early language acquisition explicit approval or disapproval by a parent of children’s utterances does not depend on whether the utterance is grammatically well formed. Furthermore, in one fascinating quantitative study it is demonstrated that the proportion of speech that is child-directed varies hugely between children in the United States and children in Yucatec Mayan villages [31]: despite the difference, children in both environments learn to speak.
The discussion of which exemplars are presented to children during learning typically focuses on the relationship between the child’s language exposure and the trajectory of the child’s language acquisition, with the suggestion that this is best studied as an example of semi-supervised learning [32, 33]. Here, however, the hope is to study how languages develop rather than how children learn, and the mixture of supervised and unsupervised exemplars in learning suggests a solution to the problem of expressivity collapse.
[…]
Fig 6: The Semi-Supervised ILM. A: An encoder, $\hat{e}$, maps $\mathcal{M}$ to $\mathcal{S}$. B: A decoder, $\hat{d}$, maps $\mathcal{S}$ back to $\mathcal{M}$. C: An autoencoder, $\hat{a}$, maps $\mathcal{M}$ to $\mathcal{M}^{\prime}$, by chaining $\hat{e}$ and $\hat{d}$. The output of each network is a vector of probabilities (denoted $p_{i}$ or $q_{i}$), which are thresholded by a function $\delta$ to deliver a binary vector, in the autoencoder, C, the signal layer is now one of three hidden layers with each node having a value between zero and one.
In the Semi-Supervised ILM, an agent is considered to have three neural networks: a decoder, $d$, an encoder, $e$, and an autoencoder, $a$, composed of $e$ and $d$:
$a=\delta\circ\hat{d}\circ\hat{e}$ (5)
or $a(m)=\delta(\hat{d}(\hat{e}(m)))$, where $\hat{e}$ is defined as the neural network mapping a meaning input to a vector of word probabilities without the final step mapping the probabilities to zeros and ones. This is illustrated in Fig. 6. During training, $\mathcal{B}$ is used in three ways: while $\mathcal{B}$ is a set of meaning-signal pairs, it is obviously possible to change the order to get signal-meaning pairs, or to ignore the signal to get a meaning. During each round of training a randomly chosen meaning-signal example from $\mathcal{B}$ is present to $e$, a randomly chosen signal-meaning is presented to $d$ and $r=20$ meanings are presented to $a$; this final training type aims to improve the ability of the autoencoder to associate the meanings with itself and is intended to mimic a child observing the world and trying to match its observations to utterances. Fig. 7 demonstrates the Semi-Supervised ILM’s ability to succeed in producing an XCS language.
A B C
Refer to caption Refer to caption Refer to caption
D E F
Refer to caption Refer to caption Refer to caption
Fig 7: The Semi-Supervised ILM evolves a stable, expressive, compositional language. The $n=8$ Semi-Supervised ILM is evolved and the performance of the language estimated. A-C plot respectively expressivity ($x$), compositionality ($c$) and stability ($s$) as a function of generation for 25 independent replicates of the $n=8$ Semi-Supervised ILM; the thick line is the mean, the thin lines show the individual replicates and the bottleneck size in all cases is 75; the autoencoder is trained using the same meanings as appear in the meaning-signal pairs used to train the decoder and encoder. In D-E $\mathcal{A}$ is selected independently to $\mathcal{B}$ and $|\mathcal{A}|=225$.
In Fig. 7(A-C) the meanings used to train the autoencoder are all meanings from $\mathcal{B}$. This may seem unrealistic: a child’s observations are not restricted to the states of the world that it has heard described. In the case of $n=8$ this does not appear to have a crucial affect on language evolution: using a independently randomly selected and larger set of meanings for autoencoder training produces similar results [Fig. 7(D-F)]. However, for larger languages this no longer holds, using a larger set, $\mathcal{A}$, of meanings selected at random and independently of $\mathcal{B}$ is required for an XCS language to evolve; this is demonstrated in Fig. 8.
A B C
Refer to caption Refer to caption Refer to caption
D
E F G
Refer to caption Refer to caption Refer to caption
Fig 8: For larger languages $\mathcal{A}$ needs to be different from, and larger than, $\mathcal{B}$. Here the $n=16$ Semi-Supervised ILM is simulated with $|\mathcal{B}|=160$. In A-C the model is run with $\mathcal{A}=\mathcal{B}$; the language does become somewhat expressive, compositional and stable, but it certainly does not evolve an XCS language. Here results for 25 replicates are shown. It may appear from these simulations that the language would continue to become more expressive, compositional and stable, but, an additional set of simulations run for 250 generations shows, D that this is not the case; here only averages are plotted, but, as before, these are based on 25 replicates. In E-G $\mathcal{A}$ is chosen independently of $\mathcal{B}$ and has size $|\mathcal{A}|=480$. Here, an XCS language quickly evolves.
In Fig. 9 a still larger language is consider: $n=20$ where the set of meanings and of signals each have over a million elements, which means that the set of expressive and compositional encoders is tiny compared to the vast number of possible maps. Since obversion would require calculating a table with $10^{12}$ entries this is too large to easily test the Obverter ILM so it is unknown whether the Obverter ILM produces an XCS language in this case. As seen in Fig. 9(A-C) For the Semi-Supervised ILM the languages become expressive and compositional but their mean stability is just a bit above $s=0.6$. This might be considered realistic and stability is a harsh measure for compositional languages, swapping just two words gives a stability of $s=0.5$. However, if this instability is a problem it is one that can be solved, for example by increasing the size of the hidden layers: Fig. 9(D-F) shows the behaviour when the hidden layer has 30 nodes, the language rapidly evolves into an XCS language.
A B C
Refer to caption Refer to caption Refer to caption
D E F
Refer to caption Refer to caption Refer to caption
Fig 9: Even larger languages do not become stable unless the hidden layer is made larger. Here $n=20$, $|\mathcal{B}|=200$ and $\mathcal{A}$ is chosen independently with $|\mathcal{A}|=600$. The usual Semi-Supervised ILM is used in A-C, with as usual 25 replicates. The language evolves some characteristics of an XCS language but does not become stable. In D-F the size of the hidden layer is increased to 30 and the language evolves extremely quickly.
A B
Refer to caption Refer to caption
Fig 10: The ideal bottleneck size also increases linearly for the Semi-Supervised ILM. Here the model is run until $x$, $c$ and $s$ exceed $\lambda_{g}=0.95$; this was done for 25 independent replicates. By testing the full relevant range of bottleneck sizes the best value of the bottleneck was estimated and this is plotted in A; the line shows the standard best linear fit, this has slope 10.5 and intercept -11.6. In every case the set of signals used to train the autoencoder is three times the size of the set of signal-meaning pairs used to train the encoder and decoder. The average number of generations to reach $\lambda_{g}=0.95$ is plotted as a function of bottleneck size in B.
The Semi-Supervised ILM demonstrates the same linear dependence on bottleneck size that was observed for the Obverter ILM, Fig. 10. The size of the bottleneck appears to be less important, or at least less finely tuned in this case, see S1 Figure.
Materials and methods
The Obverter ILM
The map $\hat{d}$ is an all-to-all feedforward neural network with one hidden layer. All layers have size $n$ and are initialized using the default flux.jl Xavier initialization. Both the hidden layer and the output layer have sigmoid non-linearities. The network is trained using stochastic gradient descent with learning rate $\eta=1.0$ for twenty epochs. In each epoch every pair in $\mathcal{B}$ is presented once; the order of presentation is random for each epoch. Since it has no tutor, the initial tutor calculates its encoder though obverting based on a decoder which is an untrained neural network.
The Semi-Supervised ILM
The maps $\hat{d}$ and $\hat{e}$ are all-to-all feedforward neural networks with one hidden layer. Except in Fig. 9(D-F) all layers have size $n$. All layers are initialized using the default flux.jl Xavier initialization. Both the hidden layer and the output layer have sigmoid non-linearities. The map $\hat{a}$ is $\hat{d}\circ\hat{e}$ and so it is an all-to-all feedforward neural network with three hidden layers. The three networks are trained at the same time using stochastic gradient descent with learning rate $\eta=5.0$. The number of epochs is again $20$. At the start of each epoch two copies of $\mathcal{B}$ are made, each in a different random order; at each iteration an item from the first copy is presented to $\hat{d}$ and an item from the other to $\hat{e}$; $r=20$ randomly chosen signals from $\mathcal{A}$ are then presented to $\hat{a}$. Having no tutor itself, the first tutor uses an untrained encoder to supervise its pupil.
Quantifying the ILM properties
Precise definitions of $x$, $c$ and $s$ are given in S2 Appendix.. These three quantities all have values between zero and one and each has a non-zero bias which depends on $n$. To make comparison across $n$ easier, in all the graphs presented here the bias has been removed so the quantity plotted is $(y-y_{0})/(1-y_{0})$ where in each case $y_{0}$ is a suitable estimate of the background value. For $x$ and $c$ this is the value calculated for a naïve agent, averaged over 40 examples, for $s$ this is the value calculated between 20 pairs of naïve agents.
Discussion
A - decoder B - encoder C - autoencoder
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fig 11: XCS languages are easier to learn. The loss function is plotted against epoch during learning for 30 generations of the Semi-Supervised ILM; the colour ranges from cold to warm as the generation number, $g$, increases. This is done separately for each of the three neural networks, the decoder A, the encoder B and the autoencoder C. In each case this is an average over 25 independent replicates. The autoencoder loss has been divided by $r=8$ to make it comparable to the other two loss values. Here $n=8$, $|\mathcal{B}|=50$ and $|\mathcal{A}|=3|\mathcal{B}|=150$.
The novel Semi-Supervised ILM introduced here demonstrates that XCS languages,languages that are expressive, compositional, and stable, arise without the use of an obverter, as a consequence of language transmission being constrained by a transmission bottleneck. While natural languages exhibit these properties to a substantial degree, they also represent a trade-off between precision, learnability and ease of production [34, 35]. That is, expressivity, compositionality and stability are not the only universal properties that are central to natural language. Fundamentally, however, in addition to being able to support communication, all languages must share one crucial property: they must be learnable.
The three properties of expressivity, compositionality and stability are central to language learnability. A stable language is, by definition, capable of being transmitted from tutor to pupil and is therefore learnable. A expressive language supports a diversity of signals and meanings and as a consequence may be challenging to learn under conditions of limited time and cognitive effort. A compositional language, however, is learnable, even if it is expressive, because a pupil need only learn the rules of composition, rather than a much larger mapping from states of the world to arbitrary, unstructured phrases.
As a simple demonstration of this, we can consider how well the neural networks inside the Semi-Supervised ILM’s agents learn as the initially arbitrary language matures into an XCS language. Of course, the learning implemented here, using stochastic gradient descent, is very different from anything that plausibly happens in the brain, so it would be wrong to suggest that the Semi-Supervised ILM’s learning dynamics align with the learning dynamics of a human language learner. However, the ease with which the neural networks reduce their loss function is likely related in some general way to how easy the language is to learn. In Fig. 11 the loss achieved by each of the pupil’s neural networks is plotted against training epoch for each generation. Since each pupil always starts in a naïve state, the initial loss is always the same at epoch one, but for higher generations, and hence more mature languages, the loss falls more quickly.
More broadly though, the issue of how well a language supports communication goes well beyond our approach here. If languages are subject to selective pressures to minimise ambiguity or reduce cognitive load on language users, these pressures might be seen to shape language change within an iterated learning model. Given its lower computational demands, the Semi-Supervised ILM could be extended to incorporate such additional pressures on language, for example related to the cognitive load experienced by a speaker while forming sentences or the ease with which these sentences can be processed or unambiguously understood by a listener.
A similar hope is that further elaboration of the model can incorporate more of the structure present in language or in putative ‘spaces of states of the world’; this includes the complex hierarchical interplay of different types of facts, be they object, both real and abstract, actions and modifiers along with a parallel set of factors related to biases, ecological salience, novelty and familiarity. Indeed, the most exciting possibility is that ILMs provide a tool for understanding how simple rules, like the generational structure and the bottleneck, can lead to more complex and more ‘formal’ rule structures emerging in languages, just as both the Obverter ILM and Semi-Supervised ILM lead to XCS languages, that is languages that incorporate apparently formal requirements like ‘a word codes for the same fact, irrespective of context’.
The ILM has a very simple tutor-pupil structure; in reality, humans do not learn from a single tutor. In some societies, such as those described in [31] or in other earlier, albeit contested, anthropological accounts such as [36], children at the age language learning occurs are embedded in a rich social environment, but are not in any obvious way ‘tutored’. Similarly the wealth of data provided by the difference in experience between children who spend time at a nursery or crèche and those who are cared for more exclusively in the family home, does not seem to produce any clear difference in the speed of language acquisition. Furthermore, language acquisition and language change continue past early childhood and, indeed, the language of teenagers is notoriously mutable. Interestingly, there is even evidence from human experiments with an artificial language that compositionality can emerge without inter-generational transmission [37]. All this indicates that a model of language development should not be restricted to isolated tutor-pupil interactions. In this, the Semi-Supervised ILM presents an advantage. In the Obverter ILM obversion takes place at a discrete transition from pupil to tutor while for the Obverter ILM there is no such transition: an agent can learn and teach at the same time. Furthermore the notion of overheard speech can be added to the model by training the signal-to-signal autoencoder and the Semi-Supervised ILM should lend itself to models where agents learn from each other as a community.
There are two obvious ways to add an autoencoder phase to learning in the ILM, the approach used here with the autoencoder mapping meaning to meaning and the ‘inside-out’ version of this with $a=e\circ d$ mapping signal to signal. The first represents the child describing observations to themselves, while the second represents the child overhearing speech and trying to guess its meaning. It is clear that both types of autoencoder could be justified as representing an aspect of the real-world situation and one approach would be to use both. For simplicity, we have not done that here, but in the tests we did do using both leads to even faster evolution of an XCS language. However, using only the signal-to-signal autoencoder does not work: it appears that one of the barriers to language evolution, at the initial stage, is that all meanings map to a small subset of the possible signals and so the set of signals generated by the tutor encoder will include only a few signals. Training the autoencoder on this small set reduces the expressivity of the language.
Currently the pupil learns in part through supervised learning. The tutor supplies input-output pairs and these are used in a classic neural network learning paradigm. However, as discussed above, directly supervised learning, as in ‘here is a cat’ / ‘what is that animal?’, is only a small part of a child’s learning experience. It seems likely that attempting to communicate and trying to understand are also important, particularly as the child becomes older. This suggests an interesting agnenet learning paradigm in which some learning occurs in a arrangement where the encoder from one agent is concatenated with the decoder from the other to form an autoencoder whose middle layer is categorical. In other words, all learning uses an autoencoder, one inside the agent and another between agents. It is not immediately obvious how to train this second autoencoder network, back-propagation cannot propagate through a categorical layer and most work-arounds for this problem involve softening the categorical step, something that is not appropriate here. Perhaps a ‘forward propagation’ approach could work; if it did it would be an interesting way to combine the ILM with a language game approach to language evolution and development.
One draw back of the ILM is that it involves a large number of meta-parameters and we have only explored as subset of these in detail. We have not presented any description of how performance depends on parameters like learning rate, the number of epochs, the number of presentations to the autoencoder in each epoch, or on any broad consideration of how details of the neural network architecture or training algorithm affect the results. Obviously we have done some preparatory exploration to find values that work, but it would be useful to examine these details more formally and more rigorously in future.
Simple models like the ILM dispense with the semiotics and logic of language; they do not allow for the possibility that an understanding of meaning and of logic might be required to account for the structure of language. The agents are also unreflective, in the sense that they do not attempt to identify deficiencies on their language with a view to improving it. This does not make them uninteresting; it is useful to note how even the simplest of models, such as the ILM, can demonstrate possible mechanisms for the presence of features in language and of the emergence of properties that might otherwise have been attributed to more complex planning.
Conclusion
The ILM is significant because it provides a mechanism for languages to develop the crucial properties of expressivity, compositionality and stability. However, while the model has been influential, it has not been used as commonly as might be expected. This may be because it uses an obverter, something that is cumbersome and unrealistic. This study introduces a novel ILM that integrates both supervised and unsupervised learning to simulate the evolution of expressive, compositional, and stable languages without relying on obversion. By employing neural networks for both encoding and decoding processes, and, crucially, combining them in the form of an autoencoder during unsupervised learning, the model can claim greater ecological validity, mirroring in this respect the complex dynamics of human language acquisition. This underscores the importance of a semi-supervised learning approach in the development of complex language structures, offering new insights into the mechanisms underlying language evolution and suggesting it may be possible to study language change in more sophisticated and larger linguistic systems.
Supporting information
S1 Figure
A B C
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
D E F
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
G H I
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Exploring optimal training set size. A-C: $\mathcal{A}=\mathcal{B}$; D-F: $|\mathcal{A}|=3|\mathcal{B}|$; G-I: $|\mathcal{B}|$ is fixed at 100. In all cases $\mathcal{A}$ and $\mathcal{B}$ are independent random samples drawn from the space of all possible signals with $n=10$ ensuring that the number of possible meaning and signals, $2^{10}$, far exceeds the largest training sets considered. Mean values for $x$, $c$ and $s$ are plotted after 20 generations (solid line) and five generations (dashed line) for 25 independent replicates.
S1 Appendix.
Obversion
Obversion, in this context, is introduced in [4] to describe an inferential step in communication akin to inversion of a signal-meaning mapping and is used in the ILM to derive an encoder map $e$ from a decoder map $d$. It is useful to introduce a mild abuse of notation. A meaning $m\in\mathcal{M}$ is a binary vector of length $n$. This is easily mapped to an integer between one and $2^{n}$, so $m$ is used here both to refer to a meaning vector and as an index for the elements of $\mathcal{M}$; the same also holds for $s$, mutatis mutandis. Now, to define obversion, a probability table is calculated:
$P_{ms}=d(s,m)$ (6)
This is the probability that the signal $s$ corresponds to the meaning $m$ according to the decoder $d$ and is calculated as described in Eq. 4. Now we define the encoder, $e$, that obverts $d$ as:
$e(m)=\text{argmax}_{s}P_{ms}$ (7)
Thus, the preferred signal $s$ corresponding to meaning $m$ is the one with the highest probability $P_{ms}$. It should be noted that $P_{ms}$ is not a probability distribution over values of $s$. While $\sum_{m}P_{ms}=1$, the same is not true for $\sum_{s}P_{ms}$.
S2 Appendix.
Quantifying language properties.
Stability measures the degree of language agreement between two language users, that is, the probability that when one agent uses their language to express a meaning $m$ using their preferred signal $s$, the other agent will be able to use their language to recover the same meaning $m$ from signal $s$.
$s=\frac{|\{m|m=d(e(m))\}|}{|M|}$ (8)
where $d$ is one agent’s decoder and $e$ is the other agent’s encoder.
Expressivity measures the degree to which a language uses different signals to express each of the possible different meanings; roughly speaking it measures the lack of ambiguity in the encoder map $e$. This is straightforward to quantify as the size of the range of $e$ as a fraction of the size of the signal set, $|S|$:
$e=\frac{|\{e(m):m\in M\}|}{|S|}.$ (9)
Compositionality quantifies the degree to which a language consistently encodes a single fact using a single word, irrespective of context. The basic idea is to quantify the extent to which this is true for each fact in turn and then average these fact-wise measures to arrive at a compositionality measure for the whole language. Recall that a fact is one component in a meaning: $m_{i}$ in $m=(m_{1},m_{2},\ldots,m_{i},\ldots,m_{n})$. For any given $m$, $m_{i}$ will always be either zero or one, and, as $m$ ranges over all its possible values, $m_{i}$ will be zero for half the time and one for the other. Now consider $s_{j}$, one of the words in $s=(s_{1},s_{2},\ldots,s_{j},\ldots,s_{n})$ where $s=e(m)$. This $s_{j}$ is also either zero or one and, indeed, considering $s=e(m)$ for all possible $m\in M$ it is possible to calculate $p(s_{j}=1|m_{i})$. Let
$h_{ij}=-p\log_{2}{p}-(1-p)\log_{2}{(1-p)}$ (10)
be the entropy for this probability distribution; a value of $h_{ij}=1$ indicates that the fact ‘1’ at meaning location $i$ maps onto the words ‘0’ and ‘1’ at signal location $j$ with equal probability, whereas an entropy value $h_{ij}=0$ indicates that the fact ‘1’ at meaning location $i$ maps exclusively to the word ‘1’ or exclusively to the word ‘0’ at signal location $j$ indicating compositionality. If $h_{i}=\text{min}_{j}(h_{ij})$ then the basic idea is that
$\tilde{c}=1-\frac{1}{n}\sum_{i}h_{i}$ (11)
However, it turns out that finding a good definition of compositionality is trickier than might be expected. While the formula for $\tilde{c}$ above does a good job for expressive languages, it does a poor job for less expressive languages. A more complicated definition is needed, though the core idea is captured by $\tilde{c}$. This more complicated definition avoids the same word being used to convey more than one fact. First, a set of optimal $h_{ij}$ is calculated, one for each fact:
$H=\{h_{ij}|h_{ij}=\text{min}_{k}(h_{ik}\text{ for }i=1\ldots n)\}$ (12)
then, for each word, the lowest value in $H$ is chosen, or, if there is no corresponding value in $H$, the word is assigned the value one:
$h^{\prime}_{i}=\left\{\begin{array}[]{ll}\text{min}_{j}h_{ij}\in H&|\{h_{ij}\in H\}|\not=0\\ 1&\text{otherwise}\end{array}\right.$ (13)
Finally,
$c=1-\frac{1}{n}\sum_{i}h^{\prime}_{i}.$ (14)
S3 Appendix.
Software and libraries
All code was written in Julia and run on Julia 1.9.3, neural networks were trained using Flux v0.14.6 and plotting used Gadfly v1.3.4 and Colors v0.12.10. Data was handled using DataFrames v1.5.0 and CSV v0.10.11. The code is available at github.com/IteratedLM/2023_12_ailm.
S4 Table.
Glossary
$n$ length of the binary meaning and signal vectors.
$\mathcal{M}$ set of meanings.
$\mathcal{S}$ set of signals.
×$\mathcal{B}\subset\mathcal{M}\times\mathcal{S}$ a set of meaning-signal pairs used for supervised learning.
$\mathcal{A}\subset\mathcal{M}$ a set of meanings used for unsupervised learning.
$\mathcal{P}$ space of $n$-vectors $(p_{1},p_{2},\ldots,p_{n})$ with $p_{i}\in[0,1]$.
meaning an element of $\mathcal{M}$ (a binary vector of length $n$)
fact one bit within a meaning
signal an element of $\mathcal{S}$ (a binary vector of length $n$)
word one bit within a signal
$d:\mathcal{S}\rightarrow\mathcal{M}$ decoder map.
$e:\mathcal{M}\rightarrow\mathcal{S}$ encoder map.
$\hat{d}:\mathcal{S}\rightarrow\mathcal{P}$ decoder neural network.
$\hat{e}:\mathcal{M}\rightarrow\mathcal{P}$ encoder neural network.
[…]
$x$ expressivity (blue in graphs).
$c$ compositionality (orange in graphs).
$s$ stability (maroon in graphs).
Acknowledgments
This work was supported by UKRI grant EP/Y028392/1 and was carried out using the computational facilities of the Advanced Computing Research Centre at the University of Bristol, www.bristol.ac.uk/acrc/. We are grateful to Dr Stewart whose philanthropy supported some of the compute resource used in this project.